OpenStudy (anonymous):

Anne walked 3.2 kilometers. How many meters did Anne walk? A. 3200 meters B. 320 meters C. 0.032 meters D. 0.0032 meters

OpenStudy (anonymous):

3200

OpenStudy (anonymous):

A

OpenStudy (anonymous):

k

OpenStudy (batman*):

Move the decimal point over three to the right. It will give you your answer. For ever 1 km its equal to 1000 meters. :) No direct answers!!!

OpenStudy (anonymous):

Jack is measuring the distance around his house. Which units should he use? A. km B. m C. cm D. mm

OpenStudy (md152727):

B

OpenStudy (anonymous):

Which equation could be used to find the perimeter of an isosceles triangle, with two sides 5 inches long, a base 6 inches long and a height of 4 inches? A. P = (6 × 4) ÷ 2 B. P = 2 (5 + 6) C. P = 5 + 6 + 4 D. P = 2 × 5 + 6

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[\huge~\rm~P=2a+b \] 

OpenStudy (anonymous):

?

pooja195 (pooja195):

We know the 2 sides are 5 in a=5 base= 6 so b=6 plug them in and you get your answer

OpenStudy (anonymous):

thanks it is d

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Which shows all the names that apply to the figure? A. parallelogram and rectangle B. parallelogram, rhombus, and rectangle C. parallelogram, rectangle, and square D. parallelogram, rhombus, rectangle, and square